# Entanglement Spectrum Dynamics as a Probe for Non-Hermitian Bulk-Boundary Correspondence in Systems with Periodic Boundaries

### v1

@@ Abstract
For the trivial to topological quench in Fig. 2(c), we observe four ES eigenvalues crossing exactly at the entanglement zero mode energy, $\xi=0.5$, as they decay. In contrast, in Fig. 2(d), we do not observe any crossings between eigenmodes, which exponentially decay towards the steady state. In the SM [35], we have numerically confirmed that the presence and absence of ES zero-crossings is determined by the phase boundary of the effective Hamiltonian $\mathsf{H}_{\rm eff}$ for OBC, given by Eq.  (12). This demonstrates a remarkable feature: the ES only cares about the topology of the OBC spectrum. We have also confirmed that the presence (absence) of ES zero-crossings is unaffected by the strength of uniform on-site balanced gain and loss (that does not affect the topology of $\hat{H}_{\rm eff}$ but do give rise to a finite lifetime to the response to a local perturbation [19, 18, 20]), consistent with our expectation that the ES zero-crossings have a topological origin [35].
@@
Figure 3(b) and (c) show the quench dynamics of ES for a PBC system, where, as before, we set the pre-quench state to be the ground state of the Hermitian SSH model in the trivial phase. In panel (b), the post-quench parameters lie in the topologically non-trivial phase for both OBC ($\omega_{\rm OBC}=1$) and PBC ($\omega_{\rm PBC}=1/2$), whereas for panel (c), quench parameters lie in the trivial phase for OBC ($\omega_{\rm OBC}=0$) and in the topological phase for PBC ($\omega_{\rm PBC}=-1/2$). We observe that two pairs of two ES eigenvalues, $\xi_{l}(t)$, crossing at $\xi=0.5$ for the trivial to topological quench shown in Fig. 3(b), while Fig. 3(c) shows that no crossings occur for trivial to trivial quenches. This demonstrates that the presence or absence of zero-crossings in the ES dynamics is determined by the OBC phase diagram of the effective Hamiltonian $\mathsf{H}_{\rm eff}$.
We briefly note that, interestingly, the ES degeneracy at the crossing point occurring in the trivial to non-trivial quench case, seen in the absence of the next-nearest neighbor hopping $J_{3}=0$ (Fig. 2(c)) and in the Hermitian case [27] is lifted when $J_{3}\neq 0$ in our open system case. We also observe that, for quenches on open chains, ES crossings remain degenerate. We do not have an explanation for these features, which remains our future work.
Conclusions.- In summary, our work has revealed that non-Hermitian topological phase transitions can be detected through the quench dynamics of the entanglement spectrum, specifically through the presence or absence of zero-crossings. Our results contribute to a deeper understanding of topological phases in open quantum systems.
@@
In this section, we show that the dynamics of the correlation function is given by Eq. (5). We also discuss the subtle difference between the non-Hermitian Hamiltonian $\hat{H}_{\rm cond}$ that governs the null-jump conditioned dynamics and the effective Hamiltonian $\hat{H}_{\rm eff}$ that governs the unconditioned dynamics of a correlation function. The GKSL equation (1) can be rewritten as,
@@
The final term is called the jump term, describing the jump process where the number of particles in the relevant system (${\rm S+\bar{S}}$) changes. In the case where one post-selects the quantum trajectories with no jump processes, the dynamics are governed by a non-Hermitian Hamiltonian $\hat{H}_{\rm cond}$.
@@
$i\partial_{t}[{\sf C}]_{j,i}(t)$ $=\sum_{a}\langle\hat{c}^{\dagger}_{i}\hat{c}_{a}\rangle\left[H_{j,a}-\frac{i}{2}(L_{j,a}+G_{j,a})\right]-\sum_{a}\langle\hat{c}^{\dagger}_{a}\hat{c}_{j}\rangle\left[H_{a,i}+\frac{i}{2}(L_{a,i}+G_{a,i})\right]+iG_{i,j}.$ (S4)
@@
Therefore, the dynamics of the correlation function are governed by $\hat{H}_{\rm eff}$. Note the sign difference in front of $\hat{G}$ compared to the conditional Hamiltonian $\hat{H}_{\rm cond}$ in Eq. (S2).
@@
We wish to obtain the relation between the spectrum of the reduced density matrix of the subsystem of interest ${\rm S}$ and the spectrum of the covariance matrix. As pointed out in the text, for free fermionic systems the reduced density matrix of the subsystem of interest is given by $\hat{\rho}_{\rm S}={\rm Tr}_{\rm\bar{S}}\hat{\rho}=e^{-\hat{H}_{\rm S}}/{\rm Tr}[e^{-\hat{H}_{\rm S}}]$. Here, $\hat{H}_{\rm S}$ is the entanglement Hamiltonian and its spectrum as the single particle ES. Since the total system Hamiltonian is Gaussian, the reduced density matrix of the subsystem of interest ${\rm S}$ should remain Gaussian. This allows us to generically express the entanglement Hamiltonian as
@@
.3 ES crossing is determined by the topology of $\hat{H}_{\rm eff}$ in OBC and is robust against local dissipation
@@
Here, we numerically confirm that the ES crossing is determined by the topology of $\hat{H}_{\rm eff}$ in OBC. Figures S1(b) and (d) show the ES dynamics after a dissipative quench in the case where the post-quench parameters are chosen to be very close to the phase boundary of the topological transition of the OBC spectrum shown in Fig. S1(a). Here, as in the situation considered in the main text, the pre-quench state is chosen to be in a topologically trivial state. As shown, the zero-crossing of the ES (does not) occur when the post-quench parameters are topological (trivial). This demonstrates that the ES crossing probes the non-Hermitian bulk-boundary correspondence in a periodic system. Furthermore, we confirm below that this property is robust against local dissipation. This is motivated by the property that was pointed out recently [19, 18, 20] that the presence of the LSE generically does not imply that correlation functions have sensitivity to boundary conditions, especially in the presence of a local dissipation [19, 18, 20]. Intuitively, this can be understood as follows: a wave packet prepared in the middle of the bulk would not know about the boundary condition unless it hits the boundary. However, as long as the wave packet has a finite lifetime, it generically decays before it propagates to the edge. A natural question to ask is whether ES zero-crossings are also sensitive to such perturbations, which we confirmed negative. Figures S1(c) and (e) show the effects of adding local gain and loss with rate $\Gamma$ described by the jump operators $\hat{L}_{A(B),j}^{{}^{\prime}}:=\sqrt{\Gamma/2}\;\hat{c}_{A(B),j}$ and $\hat{G}_{A(B),j}^{{}^{\prime}}:=\sqrt{\Gamma/2}\;\hat{c}_{A(B),j}^{\dagger}$ to the ES dynamics. As these onsite gains and losses are local, the added terms merely shift the effective Hamiltonian $\hat{H}_{\rm eff}$ in the imaginary axis and therefore do not change its topological properties. Here, we have assumed that the added loss and gain are balanced, to ensure that the system converges to an infinite temperature state, for the same reason we describe in the main text (See also Ref. [30, 37]). We see that even when we perturb the system with these series of local system-environment interactions, the property that the presence of ES zero-crossings reflects the topology of $\hat{H}_{\rm eff}$ of the post-quench parameter remain robust even near the phase boundary.
@@
Figure S1: Quench dynamics of the ES for PBC. The system is initially prepared in the trivial ground state of the SSH model  [33] $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$. Here we consider a system of 60 sites and a 10 sites subsystem S. (a) Phase diagram of the NH SSH model Eq. (11). Topological phase transitions are separated by the solid and dotted lines for OBC and PBC spectra, respectively. (b)-(e), Quench dynamics of ES for PBC. For trivial to topological quenches, the green lines represent the four eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to the zero energy entanglement value $\xi=0.5$ during the time evolution. Eigenvalues away from $\xi=0.5$ are shown in blue. (b) Trivial to topological phase quench without uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0)$. (c) Trivial to topological phase quench with uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0.1)$. (d) Trivial to trivial phase quench without uniform dissipation $(J_{1}/J_{2},\gamma/J_{2},\Gamma/J_{2})=(1.029,0.2,0)$. (e) Trivial to trivial phase quench with uniform dissipation $(J_{1}/J_{2},\gamma/J_{2},\Gamma/J_{2})=(1.029,0.2,0.1)$. All eigenmodes (blue lines) decay to $\xi=0.5$ without crossing it.

### v2

@@ Abstract
For the trivial to topological quench in Fig. 2(c), we observe four ES eigenvalues crossing exactly at the entanglement zero mode energy, $\xi=0.5$, as they decay. In contrast, in Fig. 2(d), we do not observe any crossings between eigenmodes, which exponentially decay towards the steady state. In the SM [39], we have numerically confirmed that the presence and absence of ES zero-crossings is determined by the phase boundary of the effective Hamiltonian $\mathsf{H}_{\rm eff}$ for OBC, given by Eq.  (8). This demonstrates a remarkable feature: the ES cares only about the topology of the OBC spectrum. We have also confirmed that the presence (absence) of ES zero-crossings is unaffected by the presence of disorder which respects the symmetries of the effective Hamiltonian ${\hat{H}}_{\rm eff}$ and by the strength of uniform on-site balanced gain and loss (that does not affect the topology of $\hat{H}_{\rm eff}$ but do give rise to a finite lifetime to the response to a local perturbation [2, 1, 12]), consistent with our expectation that the ES zero-crossings have a topological origin. These findings motivate us to assert that the ES crossings have a topological origin, and their presence (or absence) is determined by the topological invariants of the pre-quench Hamiltonian ${\hat{H}}$ and the effective Hamiltonian ${\hat{H}}_{\rm eff}$.
@@
Figure 3(b) and (c) show the quench dynamics of ES for a PBC system, where, as before, we set the pre-quench state to be the ground state of the Hermitian SSH model in the trivial phase. In panel (b), the post-quench parameters lie in the topologically non-trivial phase for both OBC ($\omega_{\rm OBC}=1$) and PBC ($\omega_{\rm PBC}=1/2$), whereas for panel (c), quench parameters lie in the trivial phase for OBC ($\omega_{\rm OBC}=0$) and in the topological phase for PBC ($\omega_{\rm PBC}=-1/2$). We observe that two pairs of two ES eigenvalues, $\xi_{l}(t)$, crossing at $\xi=0.5$ for the trivial to topological quench shown in Fig. 3(b), while Fig. 3(c) shows that no crossings occur for trivial to trivial quenches. This demonstrates that the presence or absence of zero-crossings in the ES dynamics is determined by the OBC phase diagram of the effective Hamiltonian $\hat{H}_{\rm eff}$. In the SM [39], we have numerically demonstrated that the findings presented in this letter hold true for non-equilibrium quenches, where the system under consideration, ${\rm S+\bar{S}}$, evolves under OBC conditions.
We briefly note that, interestingly, the ES degeneracy at the crossing point observed in the trivial to topological phase quench, which appears in the absence of next-nearest neighbor hopping ($J_{3}=0$) [Fig. 2(c)] and in the Hermitian case [30], is lifted when $J_{3}\neq 0$ in our open system scenario. Additionally, we observe that for quenches on open chains, ES crossings remain degenerate. Furthermore, the possibility that ES crossings signal a topological phase transition of the entanglement Hamiltonian, $\hat{H}_{\rm S}=-\log(Z_{\rm S}\hat{\rho}_{S})$, as reported in the quench dynamics of closed systems [30], remains an open question. We currently lack an explanation for these features, leaving this as a topic for future work.
Conclusions.— In summary, our work has revealed that non-Hermitian bulk-boundary correspondence can be detected through the quench dynamics of the entanglement spectrum, specifically through the presence or absence of zero-crossings. Our results contribute to a deeper understanding of topological phases in open quantum systems.
@@
In this section, we show that the dynamics of the two point correlation function is given by Eq. (5). We also discuss the subtle difference between the non-Hermitian Hamiltonian $\hat{H}_{\rm cond}$ that governs the null-jump conditioned dynamics and the effective Hamiltonian $\hat{H}_{\rm eff}$ that governs the unconditioned dynamics of the two point correlation function. The GKSL equation (1) can be rewritten as,
@@
The third term in Eq. (S1) is called the jump term, describing the jump process where the number of particles in the relevant system (${\rm S+\bar{S}}$) changes. In the case where one post-selects the quantum trajectories with no jump processes, the dynamics is governed by a non-Hermitian Hamiltonian $\hat{H}_{\rm cond}$.
@@
$i\partial_{t}[{\sf C}]_{j,i}(t)$ $=\sum_{a}\langle\hat{c}^{\dagger}_{i}\hat{c}_{a}\rangle\left[H_{j,a}-\frac{i}{2}(L_{j,a}+G_{j,a})\right]-\sum_{a}\langle\hat{c}^{\dagger}_{a}\hat{c}_{j}\rangle\left[H_{a,i}+\frac{i}{2}(L_{a,i}+G_{a,i})\right]+iG_{j,i}.$ (S4)
@@
Therefore, the dynamics of the two point correlation function is governed by $\hat{H}_{\rm eff}$. Note the sign difference in front of $\hat{G}$ compared to the conditional Hamiltonian $\hat{H}_{\rm cond}$ in Eq. (S2).
@@
We wish to obtain the relation between the spectrum of the reduced density matrix of the subsystem of interest ${\rm S}$ and the spectrum of the covariance matrix. As pointed out in the text, for free fermionic systems the reduced density matrix of the subsystem of interest is given by $\hat{\rho}_{\rm S}={\rm Tr}_{\rm\bar{S}}\hat{\rho}=e^{-\hat{H}_{\rm S}}/{\rm Tr}[e^{-\hat{H}_{\rm S}}]$. Here, $\hat{H}_{\rm S}$ is the entanglement Hamiltonian and its spectrum the single particle ES. Since the total system Hamiltonian is Gaussian, the reduced density matrix of the subsystem of interest ${\rm S}$ should remain Gaussian. This allows us to generically express the entanglement Hamiltonian as
@@
.3 ES crossings are determined by the topology of $\hat{H}_{\rm eff}$ under OBC and their robustness against disorder and local dissipation
Here, we numerically confirm that the ES crossings are determined by the topology of $\hat{H}_{\rm eff}$ under OBC and are robust against disorder and local dissipation. Figures S1 and S2 and show the ES dynamics after a dissipative quench under OBC and PBC, respectively. Post-quench parameters are chosen to be very close to the phase boundary of the topological transition of the OBC spectrum shown in panel (a) of Figures S1 and S2. Here, as in the situation considered in the main text, the pre-quench state is chosen to be in a topologically trivial state. As shown, zero-crossings of the ES (does not) occur when the post-quench parameters are in the topological (trivial) phase of $\hat{H}_{\rm eff}$, regardless of the boundary conditions of the system ${\rm S+\bar{S}}$. This demonstrates that the ES crossing probes the non-Hermitian bulk-boundary correspondence in a system under periodic or open boundary conditions. Table 1 summarizes the presence (or absence) of ES crossings based on the topological phase of the effective Hamiltonian, $\hat{H}_{\rm eff}$, when starting from a trivial ground state under both open and periodic boundary conditions.
System Boundary Condition Initial State Phase Effective Hamiltonian Phase ES Crossings Figure
OBC Trivial Topological Present S1(b)
OBC Trivial Trivial Absent S1(d)
PBC Trivial Topological Present S2(b)
PBC Trivial Trivial Absent S2(d)
Table 1: Presence or absence of entanglement spectrum crossings for different boundary conditions, initial phases, and quench phases.
Furthermore, we confirm below that this property is robust against local dissipation. This is motivated by the property that was pointed out recently [1, 2, 3] that the presence of the LSE generically does not imply that correlation functions have sensitivity to boundary conditions, especially in the presence of a local dissipation [1, 2]. Intuitively, this can be understood as follows: a wave packet prepared in the middle of the bulk would not know about the boundary condition unless it hits the boundary. However, as long as the wave packet has a finite lifetime, it generically decays before it propagates to the edge.
Figure S1: Quench dynamics of the ES under OBC. The system is initially prepared in the trivial ground state of the SSH model  [S37] $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$. Here we consider a system of 60 sites and a 10 sites subsystem S. (a) Phase diagram of the NH SSH model Eq. (7). Topological phase transitions are separated by the solid and dotted lines for OBC and PBC spectra, respectively. (b)-(e), Quench dynamics of ES under OBC. The green lines represent the two eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to the zero energy entanglement value $\xi=0.5$ during the time evolution. Other eigenvalues are shown in red. (b) Trivial to topological phase quench without uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0)$. A pair of ES eigenvalues cross at $\xi=0.5$ while decaying. (c) Trivial to topological phase quench with uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0.1)$. A pair of ES eigenvalues cross at $\xi=0.5$ while decaying. (d) Trivial to trivial phase quench without uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.029,0.2,0)$. All eigenmodes decay to $\xi=0.5$ without crossing it. (e) Trivial to trivial phase quench with uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.029,0.2,0.1)$. All eigenmodes decay to $\xi=0.5$ without crossing it.
Figure S2: Quench dynamics of the ES under PBC. The system is initially prepared in the trivial ground state of the SSH model  [S37] $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$. Here we consider a system of 60 sites and a 10 sites subsystem S. (a) Phase diagram of the NH SSH model Eq. (7). Topological phase transitions are separated by the solid and dotted lines for OBC and PBC spectra, respectively. (b)-(e), Quench dynamics of ES under PBC. The green lines represent the two eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to the zero energy entanglement value $\xi=0.5$ during the time evolution. Other eigenvalues are shown in blue. (b) Trivial to topological phase quench without uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0)$. A pair of ES eigenvalues cross at $\xi=0.5$ while decaying. (c) Trivial to topological phase quench with uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.009,0.2,0.1)$. A pair of ES eigenvalues cross at $\xi=0.5$ while decaying. (d) Trivial to trivial phase quench without uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.029,0.2,0)$. All eigenmodes decay to $\xi=0.5$ without crossing it. (e) Trivial to trivial phase quench with uniform dissipation $(J_{1}/J_{2},\gamma_{1}/(2J_{2}),\Gamma/(2J_{2}))=(1.029,0.2,0.1)$. All eigenmodes decay to $\xi=0.5$ without crossing it.
Figure S3: (a),(b) Quench dynamics of ES under OBC. The system is initially prepared in the trivial ground state of the SSH model $(J_{2}/J_{1},J_{3}/J_{1},\gamma_{1}/J_{1},\gamma_{2}/J_{1})=(0,0,0,0)$, with system size $L$ = 40, and a subsystem S of $M=20$ sites. At $t=0$ the relevant system $(\mathrm{S}+\bar{\mathrm{S})}$ is driven out of equilibrium by an abrupt change in parameters and allowing the system to couple to Markovian reservoirs at a rate $\gamma_{i},(i=1,2)$. The green lines represent the two eigenvalues, $\xi_{l}(t)$, of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ that are closest to $\xi=0.5$ during the time evolution. The other eigenvalues are shown in red. (a) Trivial to topological quench for OBC and PBC spectrum $(2J_{1}/(\gamma_{1}+\gamma_{2}),2J_{2}/(\gamma_{1}+\gamma_{2}),2J_{3}/(\gamma_{1}+\gamma_{2}))=(-1,1.4,1/5)$. A pair of ES eigenvalues, $\xi_{l}(t)$, cross $\xi=0.5$ as they decay. (b) Trivial to trivial (topological) quench for OBC (PBC) spectrum $(2J_{1}/(\gamma_{1}+\gamma_{2}),2J_{2}/(\gamma_{1}+\gamma_{2}),2J_{3}/(\gamma_{1}+\gamma_{2}))=(2.1,1.4,1/5)$. All the eigenvalues $\xi_{l}(t)$ decay without zero-crossings.
A natural question to ask is whether ES zero-crossings are also sensitive to such perturbations, which we confirmed negative.
Figure S4: The system is initially prepared in the trivial ground state of the SSH model under PBC for (a), and under OBC for (b), with no dissipation. Then, at $t>0$, the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ is driven out of its initial state by an abrupt change of the random hopping parameters $J_{1,j},J_{2,j}$ and a correlated dissipation with rate $\gamma_{1}$ is simultaneously turned on. The relevant system evolves according to the GKLS equation under PBC for (a), and under OBC for (b), respectively. (a), (b) ES dynamics for a subsystem S with random perturbations that preserve the symmetries of the effective Hamiltonian. (a) Quench dynamics of ES for $(\bar{J}_{1}/\bar{J}_{2},\gamma_{1}/(2\bar{J}_{2}))=(0.7,0.2)$. The system evolves under PBC. The two eigenvalues of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ closest to $\xi_{l}(t)=0.5$ are shown in green. Other eigenvalues are shown in blue. One pair of ES eigenvalues cross $\xi$ = 0.5 as they decay. (b) Quench dynamics of ES for $(\bar{J}_{1}/\bar{J}_{2},\gamma_{1}/(2\bar{J}_{2}))=(0.7,0.2)$. The system evolves under OBC. The two eigenvalues of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ closest to $\xi_{l}(t)=0.5$ are shown in green. Other eigenvalues are shown in red. One pair of ES eigenvalues cross $\xi$ = 0.5 as they decay. We set the initial state to be the (trivial) ground state at $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$, the system size $L=20$, and the size of subsystem S is $M=10$ sites. In the clean limit, $J_{1,j}=\bar{J}_{1},J_{2,j}=\bar{J}_{2}$, the quench Liouvillian lies on a topological phase for both OBC and PBC spectrum ($\omega_{\rm OBC}=1,\omega_{\rm PBC}=1$). Note crucially that the ES dynamics reflect the OBC spectrum, regardless whether the physical system is in PBC and whether the system is subject to random perturbations which preserve the symmetries of the effective Hamiltonian.
@@
Panels (c) and (e) of Figures S1 and S2 show the effects of adding local gain and loss with rate $\Gamma$ to the ES dynamics under OBC and PBC, respectively. The jump operators are given by $\hat{L}_{A(B),j}^{{}^{\prime}}:=\sqrt{\Gamma/2}\;\hat{c}_{A(B),j}$ and $\hat{G}_{A(B),j}^{{}^{\prime}}:=\sqrt{\Gamma/2}\;\hat{c}_{A(B),j}^{\dagger}$ As these onsite gains and losses are local, the added terms merely shift the effective Hamiltonian $\hat{H}_{\rm eff}$ in the imaginary axis and therefore do not change its topological properties. Here, we have assumed that the added loss and gain are balanced, to ensure that the system converges to an infinite temperature state, for the same reason we describe in the main text (See also [4, 5]). We see that even when we perturb the system with these series of local system-environment interactions, the property that the presence of ES zero-crossings reflects the topology of $\hat{H}_{\rm eff}$ remain robust even near the phase boundary. Interestingly, the presence or absence of ES crossings occurs in both PBC and OBC dynamics. However, the primary difference is that, under OBC dynamics, the time required for the first ES crossing is significantly larger compared to PBC dynamics. A similar qualitative behavior is observed in quenches where the spectrum of the effective Hamiltonian cannot be obtained through a similarity transformation ($J_{3},\gamma_{2}\neq 0$), as depicted in Figure S3.
To demonstrate that ES crossings have a topological origin, we introduced disorder to the effective Hamiltonian in two ways. First, we add disorder that preserves the sublattice symmetry as follows:
$\hat{H}_{\rm eff}^{\rm(A)}$ $=\sum\limits_{j}\Big{[}\Big{(}J_{1,j}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}+J_{2,j}\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}+J_{3,j}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j+1}\Big{)}+h.c.\Big{]}$
$+\sum_{j}\Big{(}\frac{\gamma_{1}}{2}(\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}-\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j})+\frac{\gamma_{2}}{2}(\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}-\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j})\Big{)}-i\frac{\gamma_{1}+\gamma_{2}}{2}\hat{n},$ (S10)
where $J_{\alpha,j}$ are randomly sampled from a uniform distribution over $[0.6\bar{J}_{\alpha},1.4\bar{J}_{\alpha}]$. This preserves the sublattice symmetry $\{\hat{H}_{\rm eff}^{(A)},\hat{S}\}=0$, where $\hat{S}$ satisfies $\hat{S}\hat{\bf c}(k,k^{\prime})\hat{S}^{-1}=\sigma_{z}\hat{\bf c}(-k,-k^{\prime})$, $\bm{c}(k):=(\hat{c}_{A}(k),\hat{c}_{B}(k^{\prime}))^{T}$, and we ignored the constant shift term.
Second, we introduce disorder that explicitly breaks the sublattice symmetry using a random on-site potential. The modified effective Hamiltonian in this case is given by:
$\hat{H}_{\rm eff}^{\rm(B)}$ $=\sum\limits_{j}\Big{[}\Big{(}J_{1}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}+J_{2}\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}+J_{3}\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j+1}\Big{)}+h.c.\Big{]}$
$+\sum_{j}\Big{(}\frac{\gamma_{1}}{2}(\hat{c}_{A,j}^{\dagger}\hat{c}_{B,j}-\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j})+\frac{\gamma_{2}}{2}(\hat{c}^{\dagger}_{B,j}\hat{c}_{A,j+1}-\hat{c}_{A,j+1}^{\dagger}\hat{c}_{B,j})\Big{)}-i\frac{\gamma_{1}+\gamma_{2}}{2}\hat{n}$
$+\sum_{j}W_{j}(\hat{c}_{B,j}^{\dagger}\hat{c}_{B,j}+\hat{c}_{A,j}^{\dagger}\hat{c}_{A,j}),$ (S11)
where $W_{j}$ are randomly sampled from a uniform distribution over $[-W,W]$. Sublattice symmetry is broken since $\{\hat{S},\sum_{j}W_{j}(\hat{c}_{B,j}^{\dagger}\hat{c}_{B,j}+\hat{c}_{A,j}^{\dagger}\hat{c}_{A,j})\}\neq 0$. As shown in Figure S4, our findings reveal that entanglement spectrum (ES) crossings remain robust during a quench from a trivial phase to a topological phase of the effective Hamiltonian when the disorder respects the symmetries of the effective Hamiltonian ($\hat{H}_{\rm eff}^{\rm(A)}$). In contrast, as illustrated in Figure S5 , ES crossings disappear when the disorder breaks sublattice symmetry ($\hat{H}_{\rm eff}^{\rm(B)}$).
Figure S5: The system is initially prepared in the trivial ground state of the SSH model, under PBC for (a), and under OBC for (b), with no dissipation. Then, at $t>0$, the relevant system $(\mathrm{S}+\bar{\mathrm{S}})$ is driven out of its initial state by an abrupt change of the hopping parameters $J_{1},J_{2}$, with a random on-site potential $W_{j}\in[-W,W]$, and a correlated dissipation with rate $\gamma_{1}$ is simultaneously turned on. The relevant system evolves according to the GKLS equation under PBC for (a), and under OBC for (b), respectively. (a),(b) ES dynamics for a subsystem S with random onsite disorder. We set the disorder strength to $W=0.8$. (a) Quench dynamics of ES for $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(0.5,0.1)$. The system evolves under PBC. The four eigenvalues of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ closest to $\xi_{l}(t)=0.5$ are shown in green. Other eigenvalues are shown in blue. No ES eigenvalues cross $\xi$ = 0.5 as they decay. (b) Quench dynamics of ES for $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(0.5,0.1)$. The system evolves under OBC. The four eigenvalues of the covariance matrix $\mathsf{C}_{\rm SS}(t)$ closest to $\xi_{l}(t)=0.5$ are shown in green. Other eigenvalues are shown in red. No ES eigenvalues cross $\xi$ = 0.5 as they decay. We set the initial state to be the (trivial) ground state at $(J_{1}/J_{2},\gamma_{1}/(2J_{2}))=(1.5,0)$, the system size $L=68$, and the size of subsystem S is $M=34$ sites. In the clean limit, $W_{j}=0$, the quench Liouvillian lies on a topological phase for both OBC and PBC spectrum ($\omega_{\rm OBC}=1,\omega_{\rm PBC}=1$). Note that no ES crossings are present when the system is subject to random perturbations which do not preserve the symmetries of the effective Hamiltonian.